OpenStudy (anonymous):

Instructions: Help please. Given that log 4= 0.6021, log 9= 0.9542 and log 12= 1.0792, evaluate each logarithm. Problem --> log 4.096

OpenStudy (kropot72):

Hints: \[\log_{} 4.096=\log_{} \frac{4096}{1000}=\log_{} 4096-\log_{} 1000\] \[4096=4^{6}\]

OpenStudy (anonymous):

im sorry , i dont get it . can you clarify ?

OpenStudy (kropot72):

Knowing that 4096 = 4^6 and being given the value of log 4 you can work out log 4096 by using the rule of logs: \[\log_{} a ^{b}=b \times \log_{} a\] You should know the value of log 1000.

OpenStudy (anonymous):

3 ?

OpenStudy (kropot72):

Yes. The value of log 1000 = 3. Knowing that 4096 = 4^6 and being given the value of log 4 you can work out log 4096 by using the rule of logs: \[\log_{} 4^{6}=6\times \log_{} 4=you\ can\ calculate\]

OpenStudy (anonymous):

my final answer would be 3.61, correct?

OpenStudy (kropot72):

Good work! My result is log 4.096 = 3.6126

OpenStudy (anonymous):

THANK YOU :DDD

OpenStudy (kropot72):

You're welcome :)